# Turning Sand to Gold: Recycling Data to Bridge On-Policy and Off-Policy Learning via Causal Bound

### v1

@@ B.4 Reduction from Causal to DRL Framework
The reduction we have made between the causal and the DRL framework above aligns the on-policy loss and off-policy loss terms with the corresponding terms in the causal framework. The on-policy loss is a special case of the factual loss where the observation data is filled with target treatment samples, corresponding to $\epsilon^{1}_{F}$ where $q_{1}=1$. The off-policy loss is a special case of the counterfactual loss where the observation data is fill with control treatment samples, corresponding to $\epsilon^{1}_{CF}$ where the labels and the treatment probability are shifted and $q_{0_{\text{new}}}=q_{1_{\text{old}}}=1$.
Appendix C Full Experimental Results
@@
C.1 Negligible Computational Overhead
@@
C.2 Tuning the $\lambda_{TF}$ Coefficient
In order to evaluate the impact of the $\lambda_{TF}$ coefficient on improving the loss accuracy, we conducted experiments using various $\lambda_{TF}$ values. Adding the $\lambda_{TF}$ demonstrates a consistent improvement across a range of $\lambda_{TF}$ values. This indicates that precise hyperparameter tuning is unnecessary, and as long as the SUFT OPE term is added with a coefficient within a valid range of $\lambda_{TF}$, the loss optimization improves. The $\lambda_{TF}$ value can be easily tuned through random search hyperparameter optimization within a valid range. Specifically, we encounter that a sufficient range to search from is between $\lambda_{TF}=0.5$ and $\lambda_{TF}=1.5$. The $\lambda_{TF}$ values that yield the best performance are $\lambda_{TF}=1$ for all the DQN agents and $\lambda_{TF}=5$ for the PPO agent in the Atari domain. In the MuJoCo domain, SAC performs best at $\lambda_{TF}=0.6$ and PPO at $\lambda_{TF}=1.8$.
C.3 Evaluation Formulas
To evaluate our method, we train every agent-environment configuration for 10 different runs and select the median result for each setup. Since this is an even-numbered list, there are two medians, and we chose the higher one from every run. The improvement and mean reward percentage formulas only consider environments where the numerator and denominator are greater than zero.
@@ C.3 Evaluation Formulas
(Higher reward approachRandom reward)(Lower reward approach Random reward)Lower reward approach Random $\frac{\left(\text{Higher reward approach}-\text{Random reward}\right)-\left(\text{Lower reward approach }-\text{Random reward}\right)}{\text{Lower reward approach }-\text{Random reward}}\cdot 100\%$
@@
$\log_{10}\left[\frac{\left(\text{Higher reward approach}-\text{Random reward}\right)-\left(\text{Lower reward approach }-\text{Random reward}\right)}{\text{Lower reward approach }-\text{Random reward}}\cdot 100\%+1\right]$
We add 1 to the percentage improvement to avoid the log of 0 and fractions, which result in a negative log. This addition has no impact on the results.
@@
C.4 Atari 2600 57 Games Benchmark
@@
Table 2: Training rewards comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 40 games, with statistically significant gains in 31 games. Both agents are using ${L_{1}}$ loss.
@@
Mean (%) 100.0 168.4 N/A
Figure 5: Learning curves comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using ${L_{1}}$ loss.
Figure 6: Log-scaled reward improvements comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 42 valid Atari games. SUFT outperforms the baseline agent in 29 out of the 42 valid games, achieving a reward improvement of over 100% in 26.2% of them and over 10% in 61.9% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using ${L_{1}}$ loss.
@@
Table 3: Training rewards comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 50 games, with statistically significant gains in 40 games. Both agents are using ${L_{2}}$ loss.
@@
Mean (%) 100.0 203.21 N/A
Figure 7: Learning curves comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using ${L_{2}}$ loss.
Figure 8: Log-scaled reward improvements comparison between PPO agent using the additional SUFT OPE term and the baseline agent without it across 43 valid Atari games. SUFT outperforms the baseline agent in 40 out of the 43 valid games, achieving a reward improvement of over 100% in 39.5% of them and over 10% in 83.7% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using ${L_{2}}$ loss.
Table 4: Training rewards comparison between PPO agent using the additional SUFT OPE term with different $\lambda_{TF}$ values and the baseline agent without it across 57 Atari games. All agents are using ${L_{2}}$ loss.
@@
SUFT PPO 0.1 109.46%
0.2 119.79%
0.3 127.09%
0.4 131.02%
0.5 132.93%
0.6 139.45%
0.7 149.25%
0.8 147.97%
0.9 156.59%
1.0 157.96%
1.5 172.48%
2.0 174.84%
2.5 184.85%
3.0 188.12%
5.0 203.21%
10.0 202.51%
Atari DQN Ablation Study:
Table 5: An ablation study on the DQN agent comparing the SUFT additional term impact against enlarging the buffer size, adding a Vanilla DQN, and adding a Double DQN. The DQN, Vanilla DQN, Double DQN, and SUFT DQN are using a 4K buffer size. All agents are using ${L_{2}}$ loss.
Agent Human Normalized Mean (%)
DQN 25.04%
DQN, buffer 100K 29.29%
Vanilla DQN 41.26%
Double DQN 42.31%
SUFT DQN 63.45%
This ablation study demonstrates that SUFT achieves the highest Human Normalized Mean reward gain, showing that adding the SUFT additional term is better than all the other tested components. These results highlight the potential of our method to improve the agent’s performance by enhancing sample efficiency at negligible cost, better than all other tested methods that require higher resources and computational costs. The full results for all of the agents are detailed below.
@@
Figure 10: Log-scaled reward improvements comparison between DQN agent using the additional SUFT OPE term and the baseline agent without it across 32 valid Atari games. SUFT outperforms the baseline agent in 30 out of the 32 valid games, achieving a reward improvement of over 100% in 50% of them and over 10% in 90.6% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using ${L_{2}}$ loss.
Table 7: Training rewards comparison between DQN agent using the additional SUFT OPE term with a small 4K buffer and the baseline agent without the SUFT term with a larger 100K buffer across 57 Atari games. SUFT surpasses the baseline agent in 52 games, with statistically significant gains in 47 games. Both agents are using ${L_{2}}$ loss.
@@
Figure 12: Log-scaled reward improvements comparison between Vanilla DQN agent using the additional SUFT OPE term and the baseline agent without it across 45 valid Atari games. SUFT outperforms the baseline agent in 41 out of the 45 valid games, achieving a reward improvement of over 100% in 26.7% of them and over 10% in 80% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using ${L_{2}}$ loss.
@@
Game Baseline Double DQN, buffer 4K SUFT Double DQN, buffer 4K p-value
@@
Figure 14: Log-scaled reward improvements comparison between the Double DQN agent using the additional SUFT OPE term and the baseline agent without it across 42 valid Atari games. SUFT outperforms the baseline agent in 36 out of the 42 valid games, achieving a reward improvement of over 100% in 21.4% of them and over 10% in 71.4% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using ${L_{2}}$ loss.
C.5 MuJoCo Five Environments Benchmark

### v2

@@ B.4 Reduction from Causal to DRL Framework
The reduction we have made between the causal and the DRL framework above aligns the on-policy loss and off-policy loss terms with the corresponding terms in the causal framework. The on-policy loss is a special case of the factual loss where the observation data is filled with target treatment samples, corresponding to $\epsilon^{1}_{F}$ where $q_{1}=1$. The off-policy loss is a special case of the counterfactual loss where the observation data is filled with control treatment samples, corresponding to $\epsilon^{1}_{CF}$ where the labels and the treatment probability are shifted and $q_{0_{\text{new}}}=q_{1_{\text{old}}}=1$.
Appendix C SUFT Pseudocode
C.1 SUFT DQN Pseudocode
Algorithm 1 SUFT Deep Q-learning
Initialize replay memory $D_{0}$ to capacity NN
Initialize action-value function $Q$ with random weights $\theta$
for episode =1,M=1,M do
  Initialize start state $s_{1}$
  for $t=1,T$ do
   With probability $\rho$ select a random action $a_{t}$
   otherwise select $a_{t}=\arg\max_{a}Q\left(s_{t},a;\theta_{\text{behavior}}\right)$
   Execute action $a_{t}$ in the environment and observe reward $r_{t}$ and state $s_{t+1}$
   Store transition $\left(s_{t},a_{t},r_{t},s_{t+1},Q(s_{t},a_{t};\theta_{\text{behavior}})\right)$ in $D_{0}$
   Sample random mini-batch of transitions $\left(s_{j},a_{j},r_{j},s_{j+1},Q(s_{j},a_{j};\theta_{\text{behavior}})\right)$ from $D_{0}$
   Set $y_{j}=\begin{cases}r_{j}&\text{for terminal }s_{j+1}\\ r_{j}+\gamma\max_{a^{\prime}}Q(s_{j+1},a^{\prime};\theta_{\text{target}})&\text{for non-terminal }s_{j+1}\end{cases}$
   Perform a gradient descent step with respect to the network parameters $\theta_{\text{target}}$ on:
$\hat{\mathbb{E}}_{j}\left[\left(y_{j}-Q(s_{j},a_{j};\theta_{\text{target}})\right)^{2}+\lambda_{\text{TF}}\cdot\left(Q(s_{j},a_{j};\theta_{\text{behavior}})-Q(s_{j},a_{j};\theta_{\text{target}})\right)^{2}\right]$
  end for
end for
Our method extends the standard DQN algorithm [80] in two main ways: The first is by storing the old Q values $Q(s,a;\theta_{\text{behavior}})$ in the replay buffer. The second is by incorporating the additional SUFT OPE term into the objective function $\lambda_{\text{TF}}\cdot\left(Q(s_{j},a_{j};\theta_{\text{behavior}})-Q(s_{j},a_{j};\theta_{\text{target}})\right)^{2}$, which represents the estimated treatment effect. $\theta_{\text{behavior}}$ refers to the network weights used for action selection, the behavior policy, and $\theta_{\text{target}}$ represents the current network weights, which are optimized in the gradient descent step, the target policy.
C.2 SUFT PPO Pseudocode
Algorithm 2 SUFT Proximal Policy Optimization (PPO)
Initialize a rollout buffer $D_{1}$ with horizon $T$
Initialize policy function (actor) and value function (critic) with random weights $\theta$
for iteration $1,2,\ldots$ do
  for $t=1,T$ do
   Run policy $\pi_{\theta_{\text{behavior}}}(s_{t})$
   Store transition $\left(s_{t},a_{t},r_{t},\pi_{\theta_{\text{behavior}}}(a_{t}\mid s_{t}),V(s_{t};\theta_{\text{behavior}})\right)$ in $D_{1}$
  end for
  Compute advantage estimates $\hat{A}_{1},\ldots,\hat{A}_{t}$, and rewards to go $\hat{R}_{1},\ldots,\hat{R}_{t}$
  based on the stored values $V(s_{1};\theta_{\text{behavior}}),\ldots,V(s_{t};\theta_{\text{behavior}})$
  for $k=1,K$ epochs do
   Sample random mini-batch of transitions $\left(s_{j},a_{j},r_{j},\pi_{\theta_{\text{behavior}}}(a_{j}\mid s_{j}),V(s_{j};\theta_{\text{behavior}})\right)$
   from the rollout buffer $D_{1}$
   Compute probability ratio:
$r_{j}(\theta)=\frac{\pi_{\theta_{\text{target}}}(a_{j}\mid s_{j})}{\pi_{\theta_{\text{behavior}}}(a_{j}\mid s_{j})}$
   Optimize actor loss:
$\hat{\mathbb{E}}_{j}\left[\min\left(r_{j}(\theta)\hat{A}_{j},\;\text{clip}\left(r_{j}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{j}\right)\right]$
   Optimize critic loss:
$\hat{\mathbb{E}}_{j}\left[(V(s_{j};\theta_{\text{target}})-\hat{R}_{j})^{2}+\lambda_{\text{TF}}\cdot\left(V(s_{j};\theta_{\text{behavior}})-V(s_{j};\theta_{\text{target}})\right)^{2}\right]$
  end for
end for
Our method extends the standard PPO algorithm [20] by incorporating the additional SUFT OPE term into the PPO critic loss. Notably, PPO performs multiple epochs of optimization over the same rollout buffer $D_{1}$, leading to a divergence between the behavior and target policies. This divergence is the reason PPO benefits from our method, despite being an on-policy algorithm. $\theta_{\text{behavior}}$ refers to the network weights used to interact with the environment also referred to in [20] as $\theta_{\text{old}}$, the behavior policy, and $\theta_{\text{target}}$ represents the current network weights, which are optimized, also referred to in [20] as $\theta$, the target policy. During the first epoch, before optimizing the network, the behavior and target policies are identical, so the SUFT OPE term equals zero. After the first optimization step, the target policy diverges from the behavior policy, and the SUFT OPE term now quantifies the estimated treatment effect between the target and behavior policies. In addition, the $V(s_{t};\theta_{\text{behavior}})$ values are already calculated during the calculation of the advantage estimates $\hat{A}_{t}$, and rewards to go $\hat{R}_{t}$. Thus, our method effectively reuses these stored network outputs for computing the SUFT OPE term, similar to the approach used in our SUFT DQN implementation.
Appendix D Full Experimental Results
@@
D.1 Negligible Computational Overhead
@@
D.2 Tuning the $\lambda_{TF}$ Coefficient
In order to evaluate the impact of the $\lambda_{TF}$ coefficient on improving the loss accuracy, we conducted experiments using various $\lambda_{TF}$ values. Adding the $\lambda_{TF}$ demonstrates a consistent improvement across a range of $\lambda_{TF}$ values. This indicates that precise hyperparameter tuning is unnecessary, and as long as the SUFT OPE term is added with a coefficient within a valid range of $\lambda_{TF}$, the loss optimization improves. The $\lambda_{TF}$ value can be easily tuned through random search hyperparameter optimization within a valid range. Specifically, we encounter that a sufficient range to search from is between $\lambda_{TF}=0.5$ and $\lambda_{TF}=1.5$. The $\lambda_{TF}$ values that yield the best performance are $\lambda_{TF}=1$ for all the DQN agents, $\lambda_{TF}=2.5$ for the PPO agent using the Using $L_{1}$} Loss and $\lambda_{TF}=5$ for the PPO agent using the Using $L_{2}$} Loss in the Atari domain. In the MuJoCo domain, SAC performs best at $\lambda_{TF}=0.6$ and PPO at $\lambda_{TF}=1.8$.
D.3 Evaluation Formulas
To evaluate our method, we train every agent-environment configuration for 10 different runs and report the median result for each setup. Since this is an even-numbered list, there are two central values, and we use the higher of the two as the median. The valid environments used for calculating the improvement percentage and mean reward ratio percentage formulas are those where the numerator and denominator are greater than one. This restriction allows us to exclude cases where the reward is only marginally higher than the random reward, which could otherwise yield arbitrarily large ratios. By doing so, we ensure that the metric remains stable and effectively reflects meaningful improvements.
@@ D.3 Evaluation Formulas
[…]
@@
$\log_{10}\left[\text{Improvement Percentage Formula}+1\right]$
We add 1 to the improvement percentage formula to avoid the log of 0 and fractions, which result in a negative log. This addition has no impact on the results.
@@
D.4 ˜4.1 Empirical Analysis with $L_{2}$
DRL agents are often trained using $L_{2}$ loss, for which ˜4.1 does not hold. To provide intuition for the observed $L_{2}$} performance, we empirically examine the validity of the $L_{2}$ loss for ˜4.1 using synthetic and domain-specific data. For the synthetic data, we generate one billion independent samples from the standard normal distribution, $x,y,x^{\prime},y^{\prime}\sim\mathcal{N}(0,1)$, and evaluate whether:
$\left(x-y\right)^{2}-\left(x^{\prime}-y^{\prime}\right)^{2}$ $\leq\left(x-x^{\prime}\right)^{2}+\left(y-y^{\prime}\right)^{2}$ (17)
This experiment corresponds directly to the inequality form of ˜4.1 when applied to the $L_{2}$ loss. We find that inequality 17 holds in approximately 82% of random samples, demonstrating that while ˜4.1 is not guaranteed theoretically, it is satisfied in the vast majority of cases. To further validate this observation in a domain-specific dataset, we conduct an experiment in the Atari 57 benchmark using a DQN agent trained for 400K steps with a batch size of 32. Since our method operates in an off-policy setting, the left side of the causal bound in Theorem˜4.10 is not computable from the observed data, as we don’t observe the on-policy experiences (this is the reason we optimize on the causal bound in our method). For this reason, we evaluate only inequality 11 used in the proof of Theorem˜B.12, considering the off-policy terms on both sides (with $q_{0}=1$):
$(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},\phi(x;\theta_{0}))-\mathbf{L}(y_{1},\phi(x;\theta_{1}))]]])$ $\leq$ (18)
$(\mathbb{E}_{\begin{subarray}{c}x\sim P^{0}_{X}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{0}\sim P^{x}_{Y_{0}}\end{subarray}}[\mathbb{E}_{\begin{subarray}{c}y_{1}\sim P^{x}_{Y_{1}}\end{subarray}}[\mathbf{L}(y_{0},y_{1})+\mathbf{L}(\phi(x;\theta_{0}),\phi(x;\theta_{1}))]]])$
Because the true values of $y_{0}$ and $y_{1}$ are not directly observable, since they are the true values from the given state until the end of the episode, we approximate them using the standard Bellman equation for Q-values. Under this setup, we find that inequality 18 holds in approximately 79% of cases across the 57 Atari games. These empirical findings indicate that although ˜4.1 is not theoretically guaranteed for $L_{2}$, it holds in the vast majority of the optimization steps in practice, which may explain the strong empirical performance of our method using the $L_{2}$ loss.
D.5 Empirical Evaluation of the $\delta$ Term Magnitude
In Theorem˜4.10 the $\delta$ term is ignored during optimization since it is independent of $Q$. To assess whether this omission introduces significant looseness in the causal bound, we conduct an empirical analysis on the Atari 57 benchmark using a DQN agent trained for 400K steps. Specifically, we plotted the magnitudes of the standard DQN loss, the $\delta$ term, and the SUFT OPE term throughout training. As discussed in Section˜D.4, the $\delta$ term involves the true values of $y_{0}$ and $y_{1}$, which are not directly observable. Thus, we approximate them using the standard Bellman equation for Q-values. The results, shown in Figure˜5, indicate that across all of the environments, the $\delta$ term remains consistently an order of magnitude lower than the standard DQN loss, and on the same order of magnitude as the SUFT OPE term. These findings support the decision to ignore the $\delta$ term during optimization, as it does not meaningfully loosen the causal bound.
Figure 5: Magnitude of the $\delta$ term, SUFT OPE term, and the standard DQN loss across the Atari 57 benchmark.
D.6 Atari 2600 57 Games Benchmark
@@
Table 3: Training rewards comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 40 games, with statistically significant gains in 31 games. Both agents are using $L_{1}$} loss.
@@
Mean (%) 100.0 164.7 N/A
Figure 6: Learning curves comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{1}$} loss.
Figure 7: Log-scaled reward improvements comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 41 valid Atari games. SUFT outperforms the baseline agent in 28 out of the 41 valid games, achieving a reward improvement of over 100% in 24.4% of them and over 10% in 61% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{1}$} loss.
@@
Table 4: Training rewards comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. SUFT surpasses the baseline agent in 50 games, with statistically significant gains in 40 games. Both agents are using $L_{2}$} loss.
@@
Mean (%) 100.0 197.34 N/A
Figure 8: Learning curves comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 57 Atari games. The shaded area is the 10% and 90% percentiles with linear interpolation from the 10 different seeds’ runs. Both agents are using $L_{2}$} loss.
Figure 9: Log-scaled reward improvements comparison between the PPO agent using the additional SUFT OPE term and the baseline agent without it across 42 valid Atari games. SUFT outperforms the baseline agent in 39 out of the 42 valid games, achieving a reward improvement of over 100% in 38.1% of them and over 10% in 83.3% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
Table 5: Training rewards comparison between the PPO agent using the additional SUFT OPE term with different $\lambda_{TF}$ values and the baseline agent without it across 57 Atari games. All agents are using $L_{2}$} loss.
@@
SUFT PPO 0.1 105.0%
0.2 116.05%
0.3 124.0%
0.4 126.99%
0.5 128.79%
0.6 137.06%
0.7 143.12%
0.8 146.56%
0.9 145.4%
1.0 151.72%
1.5 170.4%
2.0 167.33%
2.5 179.01%
3.0 184.66%
5.0 197.34%
10.0 198.28%
@@
Figure 11: Log-scaled reward improvements comparison between DQN agent using the additional SUFT OPE term and the baseline agent without it across 29 valid Atari games. SUFT outperforms the baseline agent in 28 out of the 29 valid games, achieving a reward improvement of over 100% in 48.3% of them and over 10% in 93.1% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
Table 7: Training rewards comparison between DQN agent using the additional SUFT OPE term with a small 4K buffer and the baseline agent without the SUFT OPE term with a larger 100K buffer across 57 Atari games. SUFT surpasses the baseline agent in 52 games, with statistically significant gains in 47 games. Both agents are using $L_{2}$} loss.
@@
Figure 13: Log-scaled reward improvements comparison between Vanilla DQN agent using the additional SUFT OPE term and the baseline agent without it across 43 valid Atari games. SUFT outperforms the baseline agent in 39 out of the 43 valid games, achieving a reward improvement of over 100% in 25.6% of them and over 10% in 79.1% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
@@
Game Baseline DDQN, buffer 4K SUFT DDQN, buffer 4K p-value
@@
Figure 15: Log-scaled reward improvements comparison between the Double DQN agent using the additional SUFT OPE term and the baseline agent without it across 40 valid Atari games. SUFT outperforms the baseline agent in 35 out of the 40 valid games, achieving a reward improvement of over 100% in 20% of them and over 10% in 72.5% of them. The red line indicates a 10% improvement, and the green line represents a 100% improvement. Both agents are using $L_{2}$} loss.
Table 11: Training rewards comparison between our causal bound optimization method and an inertia regularizer. We compare a Double DQN agent with the additional SUFT OPE term to one using an inertia regularizer across 57 Atari games. While SUFT significantly improves the baseline agent’s performance with 226.91% mean reward ratio, the inertia regularizer harms the baseline agent’s performance with only 83.21% mean reward ratio. All agents are using $L_{2}$} loss and a 4K buffer size.
Game Baseline DDQN Inertia Regularizer DDQN SUFT DDQN
Alien 773.0 762.0 806.0
Amidar 152.0 144.0 177.0
Assault 781.0 670.0 1440.0
Asterix 676.0 666.0 939.0
Asteroids 357.0 278.0 365.0
Atlantis 13900.0 3970.0 18100.0
Bank Heist 29.2 21.8 36.1
Battle Zone 10900.0 11200.0 14000.0
Beam Rider 719.0 707.0 1230.0
Berzerk 567.0 628.0 668.0
Bowling 27.8 28.8 27.0
Boxing -1.73 -3.91 18.4
Breakout 19.3 8.85 16.3
Centipede 3390.0 3840.0 3430.0
Chopper Command 1460.0 1120.0 1870.0
Crazy Climber 41200.0 42800.0 42200.0
Defender 2980.0 2650.0 4140.0
Demon Attack 369.0 191.0 2240.0
Double Dunk -23.6 -23.6 -23.4
Enduro 143.0 107.0 195.0
Fishing Derby -81.4 -84.9 -67.8
Freeway 25.2 21.2 21.3
Frostbite 344.0 273.0 265.0
Gopher 595.0 376.0 658.0
Gravitar 100.0 138.0 134.0
H.E.R.O. 4020.0 5060.0 8870.0
Ice Hockey -7.7 -6.6 -8.15
James Bond 316.0 186.0 416.0
Kangaroo 2430.0 1110.0 3280.0
Krull 3830.0 2700.0 4400.0
Kung-Fu Master 2470.0 1520.0 3680.0
Montezuma’s Revenge 0.0 0.0 0.0
Ms. Pacman 1560.0 1510.0 1650.0
Name This Game 3140.0 2420.0 4140.0
Phoenix 1080.0 640.0 2220.0
Pit Fall -14.2 -25.5 -27.5
Pong -15.9 -17.3 -15.8
Private Eye -45.2 -11.9 49.4
Q*Bert 1260.0 1180.0 1960.0
River Raid 2960.0 2550.0 3420.0
Road Runner 15700.0 11000.0 17000.0
Robotank 6.43 6.51 7.47
Seaquest 474.0 306.0 938.0
Skiing -30300.0 -26500.0 -30200.0
Solaris 1620.0 1860.0 599.0
Space Invaders 377.0 400.0 446.0
Star Gunner 970.0 980.0 1100.0
Surround -9.62 -9.37 -9.64
Tennis -23.6 -15.1 -20.3
Time Pilot 2130.0 1920.0 2640.0
Tutankham 64.1 75.8 107.0
Up N Down 3940.0 2930.0 6950.0
Venture 6.0 4.0 18.0
Video Pinball 13400.0 12400.0 15400.0
Wizard of Wor 349.0 215.0 447.0
Yars Revenge 8610.0 7990.0 10400.0
Zaxxon 4.0 12.0 10.0
Mean (%) 100.0 83.21 226.91
D.7 MuJoCo Five Environments Benchmark